International Conference on PDE & Functional and real Analysis

November 26-27, 1999 at Riso-Kaikan 2F, Science University of Tokyo See http://www.hq.sut.ac.jp/ risokai/kaikan.html for the location.

Nevember 26 (Fri.)

13:00-13:40 Soichiro Katayama (Wakayama Univ.)

Some semilinear wave equations with nonlinear damping terms

13:50-14:30 Jean-Claude Saut (Université de Paris XI) The Cauchy problem for the fifth order KP equations

15:00-15:40 Anne de Bouard (Université de Paris XI) A stochastic nonlinear Schrödinger equation

15:50-16:30 Hideo Takaoka (Tohoku Univ.)

Global wellposedness for KdV below L^{2}

November 27 (Sat.)

10:00-10:40 Nakao Hayashi (Science Univ. of Tokyo) Large time behavior of solutions to nonlinear Schrödinger equations with quadratic nonlinearities

10:50-11:30 Patrick-Nicolas Pipolo (Science Univ. of Tokyo) On solitary wave solutions to generalized KP equations 13:00-13:40 Thierry Cazenave (Université de Paris VI)

Asymptotic behavior for some conservative nonlinear wave equations

13:50-14:30 Ilia Shishmarev (Moscow State Univ.)

Asymptotic behavior for large time of solutions to the generalized Kolmogorov-Petrovskij-Piskunov equation

15:00-15:40 Tomomi Yokota (Science Univ. of Tokyo)

Monotonicity method applied to generalized complex Ginzburg-Landau equations with p-Laplacian

15:50-16:30 NGUYEN DUY Thai Son (Kyoto Sangyo Univ.)

A remark on concave-convex Hamilton-Jacobi equations

This conference is partly supported by Crant-in-Aid for Scientific Research, Ministry of Education, Japanese Government, No. 09440069 and No. 10640213.